# LightCode: Light Analytical and Neural Codes for Channels with Feedback

### v2

@@ VIII-D Interpretation of LightCode
By limiting to a symbol-by-symbol strategy, LightCode allows for better interpretability and analysis of the learned encoder representations compared to block coding schemes such as GBAF. By analyzing the power allocation and relation between encoder output and feedback from previous rounds, we draw connections between LightCode and PowerBlast.
@@ IX Conclusion
In this work, we address the problem of designing lightweight coding schemes for channels with feedback. First, we propose an analytical scheme, PowerBlast, that can be viewed as a combination of Schalkwijk-Kailath and Gallager-Nakibŏglu schemes. Using a hybrid strategy and taking advantage of the discrete nature of the signal in the final round, PowerBlast noticeably outperforms both SK and GN, providing a performance that competes with current deep learning schemes in regions of high-SNR.
@@
Appendix A Algorithm for Gallager-Nakibŏglu
@@
Round 1: Tx: Power normalization: $X_{1}=\sqrt{P_{1}}\Theta$;
Transmit: $Y_{1}=X_{1}+N_{1}$;
Rx: Send $Y_{1}$ as feedback to Tx ;
/* Tx sends the noise $U_{2}=N_{1}$ in round 2 */
Round 2: Tx: Power normalization: $X_{2}=\frac{\sqrt{P_{2}}U_{2}}{\sigma_{2}}$; where $U_{2}=N_{1}$ and $\sigma_{2}=\sigma_{ff}$;
Transmit: $Y_{2}=X_{2}+N_{2}$;
Rx: Compute the LMMSE estimate of transmit symbol $\mathbb{E}[U_{2}|Y_{2}]=\frac{\sigma_{2}\sqrt{P_{2}}Y_{2}}{1+P_{2}}$;
Update the estimate as $\hat{X_{1}}=X_{1}-\mathbb{E}[U_{2}|Y_{2}]$
/* Tx sends the error in estimate $\epsilon_{2}=\mathbb{E}[U_{2}|Y_{2}]-U_{2}$ over the next $D-2$ rounds */
while $13\leq i\leq D-$ 1 do
       Tx: Compute the error in estimate of previous round $U_{i}=\mathbb{E}[U_{i-1}|Y_{i-1}]-U_{i-1}$ ;
       Power normalization: $X_{i}=\frac{\sqrt{P_{2}}U_{i}}{\sigma_{i}}$, where $\sigma^{2}_{i}=\frac{\sigma^{2}_{i-1}}{1+S_{i-1}}$;
       Transmit: $Y_{i}=X_{i}+N_{i}$;
       Rx: Send feedback as LMMSE estimate of transmit symbol $\mathbb{E}[U_{i}|Y_{i}]=\frac{\sigma_{i}\sqrt{P_{2}}Y_{i}}{1+P_{2}}$ ;
       Update the estimate as $\hat{X_{1}}=X_{1}-\sum_{j=2}^{j=i}\mathbb{E}[U_{j}|Y_{j}]$
end while
Decoding after round $D-1$: Map $\hat{\Theta}_{D-1}=\frac{\hat{X_{1}}}{\sqrt{P_{1}}\eta}$ to the closest symbol in the $2^{K}$ PAM constellation.
/* High-SNR scheme for rounds $D$/
Round D: Tx: Compute the difference in PAM indices $U_{D}=\hat{m}-m$;
Power normalization: $X_{D}=\frac{\sqrt{P_{2}}U_{D}}{\sigma_{D}}$; where $U_{D}=\hat{m}-m$ and $\sigma_{D}=\sigma_{ff}$;
Transmit: $Y_{D}=X_{D}+N_{D}$;
Final Decoding: Use ML decoder to detect $\hat{U}$ and detect original PAM signal using $\hat{U}$.
Algorithm 4 Gallager-Nakibŏglu coding scheme

### v3

@@ VIII-D Interpretation of LightCode
By independently encoding sub-blocks of length $K=3$ and limiting to a symbol-by-symbol strategy, LightCode allows for better interpretability and analysis of the learned encoder representations compared to block coding schemes such as GBAF. By analyzing the power allocation and relation between encoder output and feedback from previous rounds, we draw connections between LightCode and PowerBlast.
@@ VIII-D2 Interpreting the encoder
We note that recent results in [32, 33] show the possibility of using post-hoc interpretability techniques to analyze the deep-learning-based codes and learn analytical approximations for the learned codes. Making progress in this direction is an interesting avenue for future work.
@@ IX Conclusion
In this work, we address the problem of designing lightweight coding schemes for channels with feedback. First, we propose an analytical scheme, PowerBlast, that can be viewed as a combination of Schalkwijk-Kailath and Gallager-Nakiboğlu schemes. Using a hybrid strategy and taking advantage of the discrete nature of the signal in the final round, PowerBlast noticeably outperforms both SK and GN, providing a performance that competes with current deep learning schemes in regions of high-SNR.
@@
Acknowledgment
This work was partly supported by ARO Award W911NF2310062, ONR Award N000142412542, NSF Award CNS-2008824, and the 6G@@@@UT center within the Wireless Networking and Communications Group (WNCG) at the University of Texas at Austin.
Appendix A Longer blocklengths for LightCode
While the input and output dimensions for LightCode can be increased to support learning longer blocklengths, converging to a good solution is harder because of the curse of dimensionality. In Table VIII, we provide the performance of LightCode trained for $K=6$ at different complexities of the encoder-decoder at forward SNR = $-1.0$ and noiseless feedback.
BLER Enc dimension Dec dimension Total # params
$3.7\times 10^{-4}$ 32 32 $7.3\times 10^{3}$
$7.2\times 10^{-5}$ 128 64 $1.09\times 10^{5}$
$2.4\times 10^{-6}$ 128 128 $4.1\times 10^{5}$
TABLE VIII: BLER for K=6 at forward SNR of $-1.0$ dB and noiseless feedback.
We begin by comparing the performance of LightCode for $K=6$ while maintaining the same encoder-decoder dimension used for $K=3$, which is $32$. From Fig. 5, the BLER of LightCode trained with $K=3$ at a forward SNR of $-1.0$ dB and noiseless feedback is $4.5\times 10^{-10}$. Hence, by transmitting $6$ bits as two sub-blocks, the BLER would be given by $1-(1-4.5\times 10^{-10})^{2}=9\times 10^{-10}$. But, as seen from Table VIII, the BLER performance for LightCode trained with $K=6$ and a hidden dimension of $32$ is orders of magnitude higher, $3.7\times 10^{-4}$. Next, we increase the hidden dimensions to $128$, increasing the total number of parameters by 100×100$\times$100 ×, and the performance is still much worse than LightCode trained with $K=3$.
Appendix B Encoder decoder architecture
Figure 12: Encoder
Figure 13: Decoder
Figure 14: Full architecture for LightCode
Appendix C Algorithm for Gallager-Nakiboğlu
1
@@ Appendix C Algorithm for Gallager-Nakiboğlu
2
3 Round 1: Tx: Power normalization: $X_{1}=\sqrt{P_{1}}\Theta$;
4 Transmit: $Y_{1}=X_{1}+N_{1}$;
5 Rx: Send $Y_{1}$ as feedback to Tx ;
6
7Round 2: Tx: Power normalization: $X_{2}=\frac{\sqrt{P_{2}}U_{2}}{\sigma_{2}}$; $U_{2}=N_{1}$ and $\sigma_{2}=\sigma_{ff}$;
8 Transmit: $Y_{2}=X_{2}+N_{2}$;
9 Rx: Compute the LMMSE estimate of transmit symbol $\mathbb{E}[U_{2}|Y_{2}]=\frac{\sigma_{2}\sqrt{P_{2}}Y_{2}}{1+P_{2}}$;
10 Update the estimate as $\hat{X_{1}}=X_{1}-\mathbb{E}[U_{2}|Y_{2}]$
11 /* Tx sends the error in estimate $\epsilon_{2}=\mathbb{E}[U_{2}|Y_{2}]-U_{2}$ over $D-2$ rounds */
12
13while $13\leq i\leq D-$ 1 do
14       Tx: Compute the error in estimate of previous round $U_{i}=\mathbb{E}[U_{i-1}|Y_{i-1}]-U_{i-1}$ ;
15       Power normalization: $X_{i}=\frac{\sqrt{P_{2}}U_{i}}{\sigma_{i}}$, $\sigma^{2}_{i}=\frac{\sigma^{2}_{i-1}}{1+S_{i-1}}$;
16       Transmit: $Y_{i}=X_{i}+N_{i}$;
17       Rx: Send feedback as LMMSE estimate of transmit symbol $\mathbb{E}[U_{i}|Y_{i}]=\frac{\sigma_{i}\sqrt{P_{2}}Y_{i}}{1+P_{2}}$ ;
18       Update the estimate as $\hat{X_{1}}=X_{1}-\sum_{j=2}^{j=i}\mathbb{E}[U_{j}|Y_{j}]$
19 Decoding after round $D-1$: Map $\hat{\Theta}_{D-1}=\frac{\hat{X_{1}}}{\sqrt{P_{1}}\eta}$ to the closest symbol in the $2^{K}$ PAM constellation.
20 /* High-SNR scheme for rounds $D$/
21
22Round D: Tx: Compute the difference in PAM indices $U=\hat{M}-M$, where $\hat{M}$ and $M$ correspond to the integer index from PAM constellation for $\hat{\Theta}$ and $\Theta$ respectively;
23 Transmit: $Y_{D}=X_{D}+N_{D}$;
24 Final Decoding: Use ML decoder to detect $\hat{U}$ and detect original PAM signal using $\hat{U}$.
Algorithm 4 Gallager-Nakiboğlu coding scheme